Assume that the apple has a mass of 0.35 kg. The raccoon throws the apple with an angle of 60 deg to the floor with an initial velocity of 4 m/s.1. Determine the maximum height that the apple will reach.
2. How long will the apple be in the air?
3. Determine the horizontal distance that the apple will travel

1
Expert's answer
2021-04-16T07:24:16-0400

1. Determine the maximum height that the apple will reach.


"h=\frac{v_y^2}{2g}=\frac{v^2\sin^2\theta}{2g}=0.612\text{ m}."

2. Find how long will the apple be in the air.


"t=2\sqrt{\frac{2h}{g}}=0.706\text{ s}."


3. Determine the horizontal distance that the apple will travel


"R=v_xt=v\cos\theta·t=1.41\text{ m}."